# Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS

### v1

@@
Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS
High-density environments, such as early galaxy overdensities, play a critical role in hierarchical structure formation and galaxy evolution, providing an ideal setting for accelerated evolution of galaxies. As part of the NIKA2 Cosmological Legacy Survey, we report the discovery of a remarkable overdensity of ten bright dusty star-forming galaxies at z5.2 in the GOODS-N field. Among these, three galaxies with spectroscopic redshifts (N2GN_1_01, 06, and 23, known as GN10, HFD850.1, and S3, respectively) had already been identified as part of the exceptional overdensity at z5.1-5.3. This structure is notably elongated along the line-of-sight, spanning approximately 30 cMpc. We report the discovery of N2GN_1_13 at zspec=5.182, a massive dusty star-forming galaxy that we confirmed as part of the overdensity through targeted NOEMA follow-up observations of the N2CLS. In addition, by combining our analysis with JWST/FRESCO data, we identified another dusty galaxy at zspec=5.201 (N2GN_1_61). In addition to these five spectroscopically confirmed galaxies, we identify five more with photometric redshifts consistent with the overdense structure. These galaxies are massive (with a median stellar mass of 1011 M⊙) and highly obscured (with a median AV of 2.9), caught in a short-lived yet extreme starburst phase at z5.2. Their high star formation rates (with a median of 680 M⊙ yr-1), efficient baryon to stellar mass conversion (20%), substantial gas reservoir and dust content, suggest rapid evolution and imminent quenching. Six of these galaxies reside in overdense filaments, while the remaining four may trace new distinct structures which will have to be spectroscopically confirmed. These few dusty galaxies dominate the star formation within the overdensity, contributing more than the numerous Hα emitters, and surpassing the cosmic average star formation rate density for this epoch. The presence of numerous very massive, dusty, and intensely star-forming galaxies at z5.2 shows that rapid stellar and dust mass assembly was already underway within the first billion years of cosmic history in overdense environments. Their properties, likely driven by efficient gas inflows along cosmic filaments in protocluster regions, suggest an accelerated evolution that current models and simulations have difficulty reproducing.
@@
At high redshift (z4), and despite recent progress, the contribution of dust-obscured galaxies to the young Universe remains quite uncertain (e.g., Traina et al. 2024). Recent ALMA surveys (as ALPINE or REBELS) showed that dust-obscured (far-IR, mm) and unobscured (UV rest-frame) star formation could have a similar contribution to the cosmic budget even at z6 (e.g., Gruppioni et al. 2020; Fudamoto et al. 2020; Khusanova et al. 2021; Barrufet et al. 2023; Fujimoto et al. 2024). In addition, recent observations by the JWST seem to show that heavily obscured sources do exist up to z12 (Rodighiero et al. 2023; Akins et al. 2023; McKinney et al. 2023; Zavala et al. 2023). These studies are not based on volume-complete surveys and/or rely on complex corrections and selections. The population of luminous and dusty galaxies at z4 is largely unconstrained and quantifying accurately the total density of the star formation rate (SFRD) at z4 from a well-controlled volume-limited survey stays an ambitious goal. One of the main contributor to the obscured SFRD at such high redshift are mm-bright, ultra massive, and optically-dark galaxies111Such galaxies are not detected in the deepest HST images (with typical 5-sigma depth of 27 mag for H band) but are detected with IRAC, and now routinely by the JWST.. A recent JWST study based on three mm-bright sources (Xiao et al. 2024) at z5-6 shows that they account for 20 % of the SFRD derived from rest-frame FUV to optical and corrected for dust attenuation at . These three galaxies are also ultra massive (log(M/M11.0) and would require, on average, about 50% of the baryons in their halos to be converted into stars. This is two to three times higher than even the most efficient galaxies at later times (Xiao et al. 2024, and references therein).
@@
The two ultra-massive galaxies in the GOODS-N field (S3 and GN10) have been found to be located in a large-scale structure in the process of formation (Herard-Demanche et al. 2025), which also contains the well known HFD850.1 galaxy. HDF850.1 is forming in one of the richest environments identified to date at z5, with 100 z=5.17-5.20 galaxies (Herard-Demanche et al. 2025; Sun et al. 2024, see also previous studies e.g. Calvi et al. 2021). Observations suggest that such (massive) overdensities are not rare at (e.g. Guaita et al. 2022; Helton et al. 2024a). They are likely progenitors of today’s massive galaxy clusters. Understanding their distribution and properties shortly after the epoch of reionization helps bridge observations of the early Universe with modern-day structures. Given the existence of a similar galaxy overdensity in both GOODS-N and GOODS-S field, Sun et al. (2024) suggest that 50%20% of the cosmic star formation at z = 5.1-5.5 occur in protocluster environments.
@@
Galaxies in high-redshift overdensities often have different properties compared to those in less dense regions. For example, Lemaux et al. (2022) found a highly significant positive relation between SFR and galaxy densities at , with galaxies in the densest proto-cluster environments exhibiting average SFR0.5 dex higher than those in the coeval field.
@@
At higher redshifts (), Li et al. (2024a) observed that galaxies in dense environments exhibit a slightly elevated SFR at a given mass compared to those in less dense regions. Additionally, they found that galaxies in high-density regions tend to have redder UV slopes, which may indicate increased dust extinction (see also Helton et al. 2024b). In the GOODS-N z5.2 overdense region, Herard-Demanche et al. (2025) compare the properties of galaxies forming within the overdensity to those outside and find that the masses, star formation rates (SFRs), and UV luminosities of galaxies inside the overdensity are significantly higher than those outside.
@@
Morishita et al. (2025) identified two overdensities of galaxies at z5.7 along the sightline of the galaxy cluster Abell 2744, where some members exhibit properties suggestive of accelerated maturation.
@@
These studies show that the environment is influencing galaxy growth early in cosmic history even if high-z galaxies are also much more gas rich than local ones, for the same total mass, which could also give rise to very high SFR.
@@
As part of the NIKA2 Cosmological Legacy Survey (N2CLS), we conducted very deep observations in GOODS-N at both 1.2 and 2 mm (Bing et al. 2023; Berta et al. 2025). We covered 160 arcmin2, to a depth close to the confusion noise of the IRAM 30 m telescope (Ponthieu et al. 2025). The large-scale overdense region at z=5.1-5.3, as observed by the FRESCO and JADES JWST programs (Herard-Demanche et al. 2025; Sun et al. 2024), is completely covered by N2CLS.
@@
The survey gives a unique opportunity to study in detail the obscured star formation in the over-dense environment and analyse the properties of its bright-mm DSFG. The paper presents in particular the discovery of N2GN_1_13 at zspec=5.182, a massive DSFG belonging to the overdensity that we characterised using NOEMA observations. Furthermore, by combining our analysis with JWST/FRESCO data, we revealed another N2CLS DSFG at zspec=5.201. We additionally identify five N2CLS galaxies with a photometric redshift of z5.3, two of which are optically dark.
The paper is organized as follows. Section 2 gives a summary of the N2CLS survey in GOODS-N. In Sect. 3, we present the NOEMA observations of N2GN_1_13 and give its properties, that we then compare with HDF850.1 and GN10 in Sect. 4. In Sect. 5 (complemented by Appendices B and D), we present the remaining 7 N2CLS DSFGs identified at z (2 with zspec, 5 with zphot) and discuss their physical properties.
Section 6 is devoted to their efficiency for converting baryonic gas into stars. In Sect 7 we report their contribution to the cosmic SFRD (complemented by Appendix C) and present their localisation in the overdense structure.
In Sect 7, we compare our results with those expected from empirical, semi-analytical and hydrodynamical models. We finally provide a summary in Sect. 9. We also discuss in Appendix A the dust production in these high-redshift galaxies.
@@
Throughout the paper, we adopt a CDM cosmology with km s-1 Mpc-1, and (Planck Collaboration et al. 2020), a Chabrier (2003) stellar initial mass function (IMF), and the Draine (2003) frequency-dependent dust absorption coefficient, , renormalized as indicated by Draine et al. (2014, i.e. m2 kg-1 as reference).
2 DSFG from the NIKA2 cosmological legacy survey in GOODS-N
@@
N2CLS is the deepest and largest survey ever made at 1.2 and 2 mm, covering COSMOS and GOODS-N (Table 1 in Bing et al. 2023). In GOODS-N, it reaches an average 1 noise level of 0.17 and 0.048 mJy over 160 sq. arcmin. Our complete sample with a purity threshold 95% comprises 65 galaxies detected with SNR in a least one of the NIKA2 bands.
@@
Thanks to a very successful NOEMA continuum observation program, combined with SMA 860 m (Cowie et al. 2017) and deep radio observations (Owen 2018), all GOODS-N sources have accurate positions that allow us to search for their multi-wavelength counterparts (Berta et al. 2025, hereafter B25). The final sample (including multiplicities) comprises 71 individual millimeter sources. Three galaxies (two identified with NOEMA, one identified with VLA) have no multi-wavelength counterpart and thus no redshift. We refer the reader to B25 for a complete description of the identification process and construction of multi-wavelength spectral energy distributions. We have spectroscopic redshifts for 29 sources222This includes the redshift measured with new NOEMA observations for N2GN_1_13 presented in this paper. (i.e. 41% of the sample) and photometric redshifts for 39 sources.
@@
Among the 71 observed galaxies, 16 are classified as HST-dark (some of which are also IRAC-dark). Two galaxies remain unidentified. For the other 14 galaxies, the average redshift is z=4.9 and the average stellar mass is M⋆= 1.91011 M⊙ (stellar masses are derived using CIGALE, see Sect 3.2). This is in agreement with the general consensus in the literature that the majority of these optically-dark galaxies are heavily dust-obscured, star-forming sources at z3-6 that dominate the high-mass end of the galaxy population at these cosmic times (e.g., Alcalde Pampliega et al. 2019; Wang et al. 2019; Barrufet et al. 2023; Xiao et al. 2023; Gottumukkala et al. 2024).
@@
We have selected one of these optically-dark N2CLS galaxies, N2GN_1_13, observed with JWST/NIRCam as part of the FRESCO program (Oesch et al. 2023) but without a confirmed redshift, as a secure target for NOEMA follow-up observations. N2GN_1_13 is a bright DSFG, with NIKA2 fluxes of 1.950.25 and 0.490.08 mJy at 1.2 and 2 mm, respectively. The accurate position of N2GN_1_13 has been secured thanks to our previous NOEMA program (W21CV, P.I. L. Bing), that detects at 2 mm a single counterpart with a flux of 39347 Jy, and a beam size of 1.61”0.93”. As illustrated in Fig. 1, the galaxy is absent in the HST images and only faintly visible in the IRAC images. The galaxy has a 6.2 single line detection in the NIRCam grism spectrum (Xiao, priv. comm.) but no conclusive redshift measurement. Assuming the NIRCam line is the Hα line, then z=5.196 and N2GN_1_13 would have the same redshift as the gigantic overdensity (Herard-Demanche et al. 2025; Sun et al. 2024).
Line | Fobs | FWHM | S | L | L |
@@ 3.1 NOEMA observations
We observed N2GN_1_13 with NOEMA in band-1 using the C-configuration as part of program W23CX (PI: G.Lagache). The total on-source integration time was 6.6 hours. The receivers were tuned to cover a frequency range suitable for detecting the CO(4-3) and CO(5-4) transitions at z=5.196.
The NOEMA data were processed using the standard calibration and reduction procedures implemented in the GILDAS333https://www.iram.fr/IRAMFR/GILDAS/ software package. This resulted in calibrated uv-tables for the lower (LSB) and upper (USB) sidebands, sampled with a spectral resolution of 20 MHz. The source 1125+596 was used as the phase and amplitude calibrator. The absolute flux uncertainty is estimated to be within 10%, while the positional accuracy is better than 0.16′′. The final dataset achieves a synthesized beam of , with a line sensitivity of 0.7 mJy/MHz and a continuum sensitivity of 5.7 Jy/beam.
We detect two spectral lines with SNR ¿ 6, which we unambiguously identify as CO(4-3) and CO(5-4) at z=5.182. As expected from the CO/[CI] flux line ratio in distant galaxies and our sensitivity, the [CI](1-0) at 79.6 GHz remains undetected with S Jy km (assuming the same FWHM as CO(4-3))444The 3 upper limit was computed as , where and are the RMS noise and the velocity width of a spectral channel, respectively.. Figure 1 presents the two CO detections, fitted along with the continuum. The best-fit Gaussian parameters derived from the line profiles are listed in Table 1.
To characterize the continuum emission, we combined all channels from the LSB and USB subbands, excluding the spectral windows containing line emission, and produced a double sideband (DSB) continuum map (Fig. 1). The map was cleaned using natural weighting with a limited number of clean components. The continuum flux was measured using two methods: integration within the 2 contour and direct measurement at the position of the emission peak. Both approaches yield a flux of 576 Jy at 84.6 GHz, suggesting that the continuum source is unresolved.

### v2

# Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star-forming galaxies at z5.2 with the N2CLS

@@
Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star-forming galaxies at z5.2 with the N2CLS
High-density environments, such as early galaxy overdensities, play a critical role in hierarchical structure formation and galaxy evolution, providing an ideal setting for accelerated galaxy growth. The GOODS-N overdensity at z5.2 has long been known, but its full extent and richness have only recently been revealed by JWST observations. It is highly elongated along the line of sight, spanning nearly 30 cMpc. We investigated its dusty galaxy population using the NIKA2 Cosmological Legacy Survey (N2CLS). Within this overdensity, we identify five luminous dusty starbursts that are spectroscopically confirmed, along with three additional candidates supported by robust photometric redshifts. Three of the spectroscopically confirmed galaxies (N2GN_1_01, 06, and 23, known as GN10, HDF850.1, and S3, respectively) had already been recognised as members of this exceptional structure. We report the discovery of N2GN_1_13 at zspec=5.182, a massive dusty star-forming galaxy that we confirmed as part of the overdensity through targeted NOEMA follow-up observations of the N2CLS. In addition, by combining our analysis with JWST/FRESCO data, we identified another dusty galaxy at zspec=5.201 (N2GN_1_61). The eight dusty galaxies are massive (with a median stellar mass of 1010 M⊙) and highly obscured (with a median AV of 3.3) and caught in a short-lived yet extreme starburst phase at z5.2. Their high star formation rates (with a median of 590 M⊙ yr-1), efficient baryon to stellar mass conversion (20% for 75% of the sample), and substantial gas reservoirs and dust content suggest rapid evolution and imminent quenching. Six of these galaxies reside in overdense filaments; the remaining two may trace new distinct structures, which will have to be spectroscopically confirmed. These few dusty galaxies dominate the star formation within the overdensity, contributing more than the numerous Hα emitters, and surpassing the cosmic average star formation rate density for this epoch. The presence of numerous very massive, dusty, and intensely star-forming galaxies at z5.2 shows that rapid stellar and dust mass assembly was already underway within the first billion years of cosmic history in overdense environments. Their properties, likely driven by efficient gas inflows along cosmic filaments in protocluster regions, suggest an accelerated evolution that current models and simulations have difficulty reproducing.
@@
At high redshifts (z4), and despite recent progress, the contribution of dust-obscured galaxies to the young Universe remains quite uncertain (e.g. Traina2024). Recent Atacama Large Millimetre/sub-millimetre Array (ALMA) surveys (such as ALPINE and REBELS) show that dust-obscured (far-IR and millimetre) and unobscured (UV rest-frame) star formation could have made a similar contribution to the cosmic budget even at z6 (e.g. Gruppioni2020; Fudamoto2020; Khusanova2021; Barrufet2023; fujimoto2024). In addition, recent JWST observations have reported candidate heavily obscured sources at z7 (Akins2023; Rodighiero2023; gandolfi25; martis2025), although spectroscopic confirmation and direct detection of the thermal dust continuum are still lacking. These studies are not based on volume-complete surveys and/or rely on complex corrections and selections. The population of luminous and dusty galaxies at z4 is largely unconstrained, and accurately quantifying the total star formation rate density (SFRD) at z4 from a well-controlled volume-limited survey remains an ambitious goal. One of the main contributors to the obscured SFRD at such high redshifts are millimetre-bright, ultra-massive, and optically dark galaxies111Such galaxies are not detected in the deepest HST images (with typical 5-sigma depth of 27 mag for H band) but are detected with IRAC, and now routinely by the JWST.. A recent JWST study based on three millimetre-bright sources (Xiao2024) at z5-6 shows that they account for 20 % of the SFRD derived from rest-frame far-UV to optical observations and corrected for dust attenuation at . These three galaxies are also ultra-massive (log(M⋆/M11.0) and would require, on average, about 50% of the baryons in their halos to be converted into stars. This is two to three times higher than even the most efficient galaxies at later times (Xiao2024, and references therein).
The two ultra-massive galaxies in the great observatories origins deep surveys north (GOODS-N) field (S3 and GN10) are located in a large-scale structure in the process of formation (Herard-Demanche2024), which also contains the well known HDF850.1 galaxy. HDF850.1 is forming in one of the richest environments identified to date at z5, with 100 z=5.17-5.20 galaxies (Herard-Demanche2024; Sun2024; see also e.g. calvi2021). Even before the extent of this structure was known, the redshifts of HDF850.1 and GN10 alone already suggested a significant overdensity of dusty star-forming galaxies at z5 in this field compared to expectations, though it remained unclear whether the individual substructures traced by these sources were physically connected (e.g. Riechers2020).
@@
Observations suggest that such (massive) overdensities are not rare at (e.g. guaita2022; helton2024). They are likely progenitors of today’s massive galaxy clusters. Understanding their distribution and properties shortly after the epoch of reionisation helps bridge observations of the early Universe with modern-day structures. Given the existence of a similar galaxy overdensity in both the GOODS-N and GOODS-S fields, Sun2024 suggest that 50%20% of the cosmic star formation at z = 5.1-5.5 occurs in protocluster environments.
@@
Galaxies in high-redshift overdensities often have different properties compared to those in less dense regions. For example, lemaux2022 find a highly significant positive relation between the star formation rate (SFR) and galaxy densities at , with galaxies in the densest protocluster environments exhibiting average SFRs 0.5 dex higher than those in the coeval field.
@@
At higher redshifts (z4.5), Li2024 observe that galaxies in dense environments exhibit a slightly elevated SFR at a given mass compared to those in less dense regions. Additionally, they find that galaxies in high-density regions tend to have redder UV slopes, which may indicate increased dust extinction (see also Helton2024_b). Herard-Demanche2024 compared the properties of galaxies forming within the GOODS-N z5.2 overdense region to those outside and found that the masses, SFRs, and UV luminosities of galaxies inside the overdensity are significantly higher than those outside.
@@
Morishita2025 identify two overdensities of galaxies at z5.7 along the sightline of the galaxy cluster Abell 2744, where some members exhibit properties suggestive of accelerated maturation.
@@
These studies show that the environment influences galaxy growth early in cosmic history even if high-z galaxies are also much more gas-rich than local ones, for the same total mass, which could also give rise to very high SFRs.
@@
As part of the NIKA2 Cosmological Legacy Survey (N2CLS), we conducted very deep observations of GOODS-N at both 1.2 and 2 mm (Bing2023; berta2025; bethermin2025). We covered 160 arcmin2 and to a depth close to the confusion noise of the Institut de Radioastronomie Millimétrique (IRAM) 30 m telescope (Ponthieu2025). The large-scale overdense region at z=5.1-5.3, as observed by the FRESCO and JADES JWST programmes (Herard-Demanche2024; Sun2024), is completely covered by N2CLS.
@@
The survey provides a unique opportunity to study in detail the obscured star formation in the over-dense environment and analyse the properties of its millimetre-bright dusty star-forming galaxy (DSFG). This paper presents the discovery of N2GN_1_13 at zspec=5.182, a massive DSFG belonging to the overdensity that we characterised using Northern Extended Millimetre Array (NOEMA) observations. Furthermore, by combining our analysis with JWST/FRESCO data, we revealed another N2CLS DSFG at zspec=5.201. We additionally identify three N2CLS galaxies with a photometric redshift of z5.3, one of which is optically dark.
The paper is organised as follows. Section 2 gives a summary of the N2CLS survey of GOODS-N. In Sect. 3 we present the NOEMA observations of N2GN_1_13 and give its properties, which we then compare with those of HDF850.1 and GN10 in Sect. 4. In Sect. 5 (complemented by Appendices B and C) we present the remaining five N2CLS DSFGs identified at z (two with zspec, three with zphot) and discuss their physical properties. In Sect. 6 we discuss the efficiency of the DSFGs in converting baryonic gas into stars. In Sect. 7 we report their contribution to the cosmic SFRD and present their localisation in the overdense structure. In Sect. 8 we compare our results with those expected from empirical, semi-analytical, and hydrodynamical models. We finally provide a summary in Sect. 9. We also examine the dust production in these high-redshift galaxies in Appendix A, and present another N2CLS galaxy candidate at z5.2 (N2GN_1_20) in Appendix D.
@@
Throughout the paper, we adopt a cold dark matter cosmology with km s-1 Mpc-1, and (planck2020_cosmo), a chabrier2003 stellar initial mass function (IMF), and the draine2003 frequency-dependent dust absorption coefficient, , re-normalised as indicated by draine2014.
2 DSFGs from the NIKA2 cosmological legacy survey in GOODS-N
@@
N2CLS is the deepest and largest survey ever made at 1.2 and 2 mm, covering COSMOS and GOODS-N (Table 1 in Bing2023). In GOODS-N, it reaches an average 1 noise level of 0.17 and 0.048 mJy over 160 sq. arcmin. Our complete sample with a purity threshold 95% comprises 65 galaxies detected with S/N in a least one of the NIKA2 bands.
@@
Thanks to a very successful NOEMA continuum observation programme, combined with sub-millimetre array (SMA) 860 m (Cowie2017) and deep radio observations (owen2018), all GOODS-N sources have accurate positions that allow us to search for their multi-wavelength counterparts (berta2025, hereafter B25). The final sample (including multiplicities) comprises 71 individual millimetre sources. Three galaxies (two identified with NOEMA, one identified with the VLA) have no multi-wavelength counterpart and thus no redshift. We refer the reader to B25 for a complete description of the identification process and construction of multi-wavelength spectral energy distributions (SEDs). The paper also provides a full description of the data employed for the SED fitting, which have been released and are publicly accessible222https://data.lam.fr/n2cls. We have spectroscopic redshifts for 29 sources333This includes the redshifts measured with new NOEMA and JWST observations for N2GN_1_13 and N2GN_1_61 presented in this paper. (i.e. 41% of the sample) and photometric redshifts for 39 sources.
@@
Among the 71 individual galaxies, 16 are classified as Hubble Space Telescope (HST)-dark (some of which are also IRAC-dark). Two HST-dark galaxies remain unidentified despite their NOEMA (N2GN_1_17_b) and Very Large Array (VLA; N2GN_1_55) counterpart. For the other 14 galaxies (6 with spectroscopic redshifts and 8 with photometric redshifts), the average redshift is z=4.9 and the average stellar mass is M⋆= 3.91011 M⊙ (stellar masses are derived using CIGALE; see Sect 3.2). This is in agreement with the general consensus in the literature that the majority of these optically dark galaxies are heavily dust-obscured, star-forming sources at z3-6 that dominate the high-mass end of the galaxy population at these cosmic times (e.g. alcade2019; wang2019; Barrufet2023; xiao2023; Gottumukkala2024).
@@
We selected one of these optically dark N2CLS galaxies, N2GN_1_13, observed with JWST/NIRCam as part of the FRESCO programme (Oesch2023) but without a confirmed redshift, as a secure target for NOEMA follow-up observations. N2GN_1_13 is a bright DSFG, with NIKA2 fluxes of 1.950.25 and 0.490.08 mJy at 1.2 and 2 mm, respectively. The accurate position of N2GN_1_13 has been secured thanks to our previous NOEMA programme (W21CV, P.I. L. Bing), which detected at 2 mm a single counterpart with a flux of 39347 Jy and a beam size of 1.61”0.93”. As illustrated in Fig. 1 (see also Fig. 7), the galaxy is absent in the HST images and emerges in JWST images at 3.3 m. The galaxy has a 6.2 single line detection in the NIRCam grism spectrum (Xiao, priv. comm.) but no conclusive redshift measurement. Assuming the NIRCam line is the Hα line, then z5.196 and N2GN_1_13 would have the same redshift as the gigantic overdensity (Herard-Demanche2024; Sun2024).
| Line | FWHM | S | L | L | |
@@ 3.1 NOEMA observations
We observed N2GN_1_13 with NOEMA in band-1 using the C-configuration as part of programme W23CX (PI: G.Lagache). The total on-source integration time was 6.6 hours. The receivers were tuned to cover a frequency range suitable for detecting the CO(4-3) and CO(5-4) transitions at z5.196.
The NOEMA data were processed using the standard calibration and reduction procedures implemented in the GILDAS555https://www.iram.fr/IRAMFR/GILDAS/ software package. This resulted in calibrated uv tables for the lower and upper sidebands, sampled with a spectral resolution of 20 MHz. The source 1125+596 was used as the phase and amplitude calibrator. The absolute flux uncertainty is estimated to be within 10%, while the positional accuracy is better than 0.16′′. The final dataset achieves a synthesised beam of , with a line sensitivity of 0.7 mJy/MHz and a continuum sensitivity of 5.7 Jy/beam.
We detect two spectral lines with S/N ¿ 6, which we unambiguously identify as CO(4-3) and CO(5-4) at z=5.182. As expected from the CO/[CI] flux line ratio in distant galaxies and our sensitivity, the [CI](1-0) at 79.6 GHz remains undetected with S Jy km (assuming the same full width at half maximum as CO(4-3))666The 3 upper limit was computed as , where and are the RMS noise and the velocity width of a spectral channel, respectively.. Figure 1 presents the two CO detections, fitted along with the continuum. The best-fit Gaussian parameters derived from the line profiles are listed in Table 1.
To characterise the continuum emission, we combined all channels from the lower and upper sidebands, excluding the spectral windows containing line emission, and produced a double sideband continuum map (Fig. 1). The map was cleaned using natural weighting with a limited number of clean components. The continuum flux was measured using two methods: integration within the 2 contour and direct measurement at the position of the emission peak. Both approaches yield a flux of 576 Jy at 84.6 GHz, suggesting that the continuum source is unresolved.